4. A runner maintains a constant speed of 6.0 m/s with a stride length of 1.4 m. If the runner
increases his stride length to 1.5 m without changing his stride frequency, what will his new speed
be?
1
Expert's answer
2019-11-04T10:39:25-0500

Frequency (number of steps per second): "f= {v} \/ {l}=6.0 \/ 1.4 = 4.286 sec^{-1}"

If the step length increases : "v= f * l = 4.286 * 1.5 = 6.43 meter\/sec"